Items bought by for R158 are sold for R120. The loss expressed as a percentage of cost price is?


1
Expert's answer
2021-06-12T00:23:21-0400

The loss expressed as a percentage of cost price is "\frac{158-120}{158}100\%=\frac{38}{158}100\%\approx 24\%."